My teacher gave us a worksheet and one of the problems said:
"The answers to an absolute value problem are graphed below. What was the equation?"
The graph had -12 and 120 graphed. I know that I find the medain number between the two, but then what should I do?

Answer by stanbon(75887)

"The answers to an absolute value problem are graphed below. What was the equation?"
The graph had -12 and 120 graphed. I know that I find the medain number between the two, but then what should I do?
---------------------------
The median number is (-12+120)/2 = 108/2 = 54
----------
The distance from the median to an end point is 120-54 = 66
------------
Absolute value is an expression of distance:
|x-54| = 66
This means "The distance each of your numbers is from 54 is 66.
One of the numbers is 66 above 54.
The other number is 66 below 54.
================
Cheers,
Stan H.